# The Polar Express: Optimal Matrix Sign Methods and Their Application to the Muon Algorithm

### v3

@@ Abstract
Computing the polar decomposition and the related matrix sign function has been a well-studied problem in numerical analysis for decades. Recently, it has emerged as an important subroutine within the Muon algorithm for training deep neural networks. However, the requirements of this application differ sharply from classical settings: deep learning demands GPU-friendly algorithms that prioritize high throughput over high precision. We introduce Polar Express, a new method for computing the polar decomposition. [^1] https://github.com/NoahAmsel/PolarExpress Like Newton–Schulz and other classical polynomial methods, our approach uses only matrix-matrix multiplications, making it very efficient on GPUs. Inspired by earlier work of Chen & Chow and Nakatsukasa & Freund, Polar Express adapts the update rule at each iteration by solving a minimax optimization problem. We prove that this strategy minimizes error in a worst-case sense, allowing Polar Express to converge as rapidly as possible both in the early iterations and asymptotically. We also address finite-precision issues, making it practical to use in bfloat16. When integrated into the Muon training framework, our method leads to consistent improvements in validation loss when training a GPT-2 model on one billion tokens from the FineWeb dataset, outperforming recent alternatives across a range of learning rates.
@@ 1  Introduction
Advanced linear algebra is making its way into deep learning. Efficient algorithms for computing matrix functions have found exciting new applications in training neural networks. In particular, approximations to the matrix-inverse are used in the full Adagrad method [15], the matrix square-root and quarter-root appear as subroutines in the Shampoo and Soap optimizers [20, 48, 51], and most recently, the matrix sign function has become a key ingredient of the Muon optimizer [5, 4, 25].
@@
While the problem of computing these matrix functions has been studied by numerical analysts for decades, applications in deep learning come with different requirements than those in computational science. For deep learning, it is critical to take maximum advantage of GPU-friendly operations like matrix-matrix products and to avoid less parallel operations. Moreover, memory overhead must be small to handle large models. On the other hand, high accuracy is typically less important; the gold standard of sixteen digits of accuracy is overkill in deep learning.

### v4

@@ Abstract
Computing the polar decomposition and the related matrix sign function has been a well-studied problem in numerical analysis for decades. Recently, it has emerged as an important subroutine within the Muon optimizer for training deep neural networks. However, the requirements of this application differ sharply from classical settings: deep learning demands GPU-friendly algorithms that prioritize high throughput over high precision. We introduce Polar Express, a new method for computing the polar decomposition. [^1] https://github.com/NoahAmsel/PolarExpress Like Newton-Schulz and other classical polynomial methods, our approach uses only matrix-matrix multiplications, making it very efficient on GPUs. Inspired by earlier work of Chen & Chow and Nakatsukasa & Freund, Polar Express adapts the update rule at each iteration by solving a minimax optimization problem. We prove that this strategy minimizes error in a worst-case sense, allowing Polar Express to converge as rapidly as possible both in the early iterations and asymptotically. We also address finite-precision issues, making it practical to use in bfloat16. When integrated into Muon, our method yields consistent improvements in validation loss for a GPT-2 model trained on one to ten billion tokens from the FineWeb dataset, outperforming recent alternatives across a range of learning rates.
@@ 1 Introduction
Advanced linear algebra is making its way into deep learning. Efficient algorithms for computing matrix functions have found exciting new applications in training neural networks. In particular, approximations to the matrix-inverse are used in the full Adagrad method (Duchi et al., 2011), the matrix square-root and quarter-root appear as subroutines in the Shampoo and Soap optimizers (Gupta et al., 2018; Shi et al., 2023; Vyas et al., 2025), and most recently, the matrix sign function has become a key ingredient of the Muon optimizer (Bernstein and Newhouse, 2024b; a; Jordan et al., 2024b). While the problem of computing these matrix functions has been studied by numerical analysts for decades, applications in deep learning come with different requirements than those in computational science. For deep learning, it is critical to take maximum advantage of GPU-friendly operations like matrix-matrix products and to avoid less parallel operations. Moreover, memory overhead must be small to handle large models. On the other hand, high accuracy is typically less important; the gold standard of sixteen digits of accuracy is overkill in deep learning.